# Cautionary tales on heating-rate prescriptions in kilonovae

### v1

@@ 3 Bias due to misspecified prescription
As we alluded to above, a critical question to investigate is what is the impact of using the incorrect prescription for the heating rate given observational uncertainties. This is pertinent as if the impact is smaller than observational uncertainties, then pursuing a more detailed prescription is largely an academic exercise. The impact shown in the lightcurves in Sec. 2 suggests a significant difference, here, we aim to explore what the impact translates to in terms of fitting observations.
@@
We simulate target-of-opportunity style observations with the Vera Rubin Observatory (Ivezić et al., 2019) for a kilonova with the updated heating-rate prescription using Redback. We then fit the simulated data with a two component kilonova model with the old heating-rate prescription (Korobkin et al., 2012), i.e., a misspecified model and the updated prescription using Redback with the pymultinest (Buchner, 2016) sampler implemented in bilby (Ashton et al., 2019). In Fig. 3, we show this data and fits with a misspecified model (in the left panel) and the correct model (right panel). These fits highlight a key point, by eye, even an incorrect prescription can provide great fits to the observations, that would not lead to any suspicion that something is wrong with the models. We note that methods such as Bayesian model selection could elucidate the correct model with a $\ln\rm{BF}\approx 40$, i.e., a $e^{40}$ preference for the updated, (input) heating-rate prescription. However, Bayesian model selection is often not practical outside of simulated data sets where we can control all aspects of data generation.
@@
While the fits with both prescriptions look good and do not raise suspicion, the same is not true of the parameter recovery. In Fig. 4, we show violin plots of the posteriors of the estimated ejecta masses and velocities of the two components, with the fit with the updated heating-rate prescription in blue and the older prescription in orange. The black crosses indicate the true input parameters. As illustrated, the incorrect prescription, while providing good fits by eye to the observations, fail to recover the true input parameters; estimating an ejecta mass of ${0.006}_{-0.001}^{+0.001}$ and ${0.069}_{-0.007}^{+0.006}~{}M_{\odot}$ ($68\%$ credible interval) for the two components respectively, fully excluding the true input parameters of $0.02$ and $0.025~{}M_{\odot}$, respectively. Meanwhile, the velocities, while correctly recovered, are significantly more unconstrained by the older prescription relative to the updated heating-rate prescription, with a $68\%$ credible-interval width of $11\%$ relative to $0.5\%$ with the updated prescription. The unconstrained posterior is a consequence of the older prescription providing no diagnostic power towards the velocity from the shape of the lightcurve.
@@
Figure 5: Recovered posteriors for salient parameters from a fit with no underlying error on the heating-rate (orange) and one with a $10\%$ uncertainty (blue). The black lines indicate the true input of the simulated data.
@@ 4 Uncertainty in heating rate prescriptions
For example, for low electron fractions, $Y_{e}\lesssim 0.2$, the choice between the Duflo-Zuker (Duflo & Zuker, 1995) and the Finite Range Droplet Model (FRDM) mass formulae (Möller et al., 1995) produce a difference of up to an order of magnitude in the heating rate (Barnes et al., 2016; Rosswog et al., 2017). To investigate the effect of this uncertainty, we simulate synthetic kilonova observations, with both components with low electron fractions of $0.01$ and $0.05$, respectively, consistent with the ejecta of a neutron-star black hole merger. We fit these observations but increase the heating rate for low electron fractions by a factor of $5$. The impact of a discrepancy in heating rate up to a factor of $5$ is severe; we fail to recover the correct input masses and velocities, with no posterior overlap with the true input. In particular, as the heating rate is assumed to be larger (i.e., more energy released per gram of material), we infer significantly smaller ejecta masses; ${0.003}_{-0.002}^{+0.001}$ and ${0.004}_{-0.001}^{+0.001}~{}M_{\odot}$ cf. the true input masses $0.015$ and $0.03~{}M_{\odot}$. The large difference in heating rate also translates into a visually discernible and significantly worse quality of fit. This large difference in fit quality and potential additional diagnostics from spectral observations suggest that data alone could distinguish between the heating-rate predictions for the two different mass formulae. However, if the effect of the mass formula is smaller, e.g., up to a factor of $\approx 2$, then photometric data alone is likely not be sufficient to distinguish between the two mass formulae, with no discernible difference in the quality of the fit. For this simulated study we recover component masses of ${0.009}_{-0.001}^{+0.001}$ and ${0.046}_{-0.020}^{+0.004}~{}M_{\odot}$, with no posterior overlap with the true input masses mentioned above and velocities of ${0.227}_{-0.004}^{+0.003}$ and ${0.016}_{-0.004}^{+0.004}$ c (cf. true input of $0.25$ and $0.1$ c), again, with no overlap with the true input. These discrepancies are concerning, as if the discrepancy in the heating-rate due to the different mass formulae is not substantially large ($\gtrsim 3$), such that data cannot readily distinguish between the two models, then estimates of the mass of velocity for lanthanide-rich kilonova ejecta are at best limited to a precision of $\approx 50\%$. This is significantly larger than the statistical uncertainties and far larger than the precision required for any precision constraints on the chemical enrichment of the Universe, or multi-messenger constraints on the equation of state or Hubble constant.
Other sources of uncertainty in the heating rate, such as due to the nuclear network, reaction rates and thermalization are less easy to quantify. However, it is reasonable to assume that it is $\lesssim 10\%$, given the range of estimates from multiple groups (Barnes et al., 2016; Hotokezaka & Nakar, 2020). We now explore the systematic uncertainty on inferences into the properties of the ejecta given the underlying uncertainty of $10\%$ on the heating rate. We again simulate observations for a kilonova with the updated heating-rate prescription. However, we now fit the observations by adding an extra parameter, $f_{\rm fudge}$, which perturbs each term in the heating rate prescription, i.e., Eq. 2. We set a uniform prior on this parameter from $0.95-1.05$. This is, in essence, marginalizing our inferences accounting for a $10\%$ error on each term in the heating-rate prescription. This is likely conservative, as we are perturbing all terms by the same parameter rather than allowing perturbations on each term to vary.
We show posteriors for salient parameters obtained by this result with those where $f_{\rm fudge}=1$ i.e., no heating rate uncertainty in Fig. 5. As expected, our posteriors marginalizing over the heating-rate uncertainty are broader. In particular, our precision at $90\%$ credible interval on the velocities decreases from $10$ and $22\%$ for each component, to $15$ and $35\%$ when marginalising over the heating-rate uncertainty. For the ejecta masses, we see an $\approx 5\%$ decrease in precision. The overall increase in size of the credible interval width can be interpreted as a systematic uncertainty in our estimated properties due to $10\%$ uncertainty on the heating-rate prescription. In other words, given our estimated uncertainty on the heating rate of kilonovae, the absolute precision of estimated ejecta masses and velocity is $\approx 5$ and $\approx 10\%$, respectively, significantly larger than the statistical uncertainties from analyses on AT2017gfo. These systematic uncertainties limit the precision and use of inferred properties of kilonovae to understand fundamental physics such as multi-messenger constraints on the equation of state or estimates of the Hubble constant, their population properties, and their role in the chemical enrichment of the Universe.
@@ 5 Conclusions
In this Letter, we have highlighted the dangers of using kilonova heating-rate prescription that are based on a very narrow $Y_{e}$ range such as the one from Korobkin et al. (2012). Although this prescription has been widely used as a key ingredient for estimating kilonova properties and designing survey strategies, it is unsuitable for electron fractions other than $Y_{e}=0.04$. Here, we have shown the differences in individual lightcurves and populations from an updated prescription following Rosswog & Korobkin (2024). We make these lightcurve models and heating-rate prescription publicly available in open-source package, Redback (Sarin et al., 2023).
@@
We also show the dangers of using the incorrect prescription by fitting realistic, simulated observations with the Vera Rubin Observatory. While fits look great by eye and do not raise suspicion, they fail to recover the correct input ejecta masses at greater than $5\sigma$ and significantly under-constrain the velocities. This has disastrous consequences for judging the relevance of neutron star mergers for the chemical enrichment of the cosmos, the interpretation of kilonovae and their use in multi-messenger studies to constrain the equation of state or Hubble constant. In a real observation, where we do not know the true ‘input’ parameters, this highlights the need to verify estimated parameters with physical intuition.
@@
Our study demonstrates that a conservative estimate of the uncertainty in heating rates translates in a systematic uncertainty of $\approx 5$ and $\approx 10\%$ on the ejecta masses and velocity, respectively, for lanthanide-poor ejecta. For lanthanide-rich ejecta, the systematic uncertainty is larger up to $\approx 50\%$ due to uncertainties in the nuclear mass formulae. This systematic uncertainty limits the precision of constraints on equation of state and Hubble constant from kilonovae, and should ideally be marginalised over when fitting observations such as AT2017gfo, or as additional components in gamma-ray burst afterglows, e.g., GRB 211211A and GRB230307A to ensure that the interpretation of these events are not affected. Moreover, this raises the need for caution regarding precision constraints based on kilonovae properties and interpretations of their role in the chemical enrichment of the Universe.
@@ 6 Acknowledgments
We are grateful to Oleg Korokbin for helpful comments on the manuscript. N.S. is supported by a Nordita Fellowship. Nordita is supported in part by NordForsk. S.R. has been supported by the Swedish Research Council (VR) under grant number 2020-05044, by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012, by the Knut and Alice Wallenberg Foundation under grant Dnr. KAW 2019.0112, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and by the European Research Council (ERC) Advanced Grant INSPIRATION under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101053985).

### v2

@@ 3 Bias due to misspecified prescription
As we alluded to above, a critical question to investigate is what is the impact of using the incorrect prescription for the heating rate given observational uncertainties. This is pertinent as if the impact is smaller than observational uncertainties, then pursuing a more detailed prescription is largely an academic exercise. The impact shown in the lightcurves in Sec. 2 suggests a significant difference, here, we aim to explore what the impact translates to in terms of fitting observations. We note while we only apply this methodology here to exploring the impact of differences in “naked” heating-rate prescriptions, such a procedure is equally valid to explore the impact of other systematics in kilonova modelling.
We simulate target-of-opportunity style observations with the Vera Rubin Observatory (Ivezić et al., 2019) for a kilonova with the updated heating-rate prescription using Redback. In particular, we set the input parameters for the simulated data to $0.015$ and $0.03~{}M_{\odot}$ for the ejecta mass of the two components, with floor temperatures of $3000$ and $1200$ K. We set the velocities to $0.25$ and $0.1$c and the electron fractions to $0.4$ and $0.15$. We then fit the simulated data with a two component kilonova model with the old heating-rate prescription (Korobkin et al., 2012), i.e., a misspecified model and the updated prescription using Redback with the pymultinest (Buchner, 2016) sampler implemented in bilby (Ashton et al., 2019). We note that we fit for all $8$ model parameters i.e., the ejecta mass, velocity, floor temperature and electron fractions for each component. We also place two constraints on our prior, such that the ejecta mass of the second component (likely the disk-wind) is larger but moving at a slower velocity than the dynamical ejecta, consistent with expectations from numerical simulations (e.g., Perego et al., 2014; Dietrich et al., 2017; Siegel & Metzger, 2018).
@@
In Fig. 3, we show this data and fits with a misspecified model (in the left panel) and the correct model (right panel). These fits highlight a key point, by eye, even an incorrect prescription can provide great fits to the observations, that would not lead to any suspicion that something is wrong with the models. This highlights the need to check for consistencies from inferences on lightcurves with other observables such as overall constraints on energetics or the spectrum. We note that methods such as Bayesian model selection could elucidate the correct model with a $\ln\rm{BF}\approx 20$, i.e., a $e^{20}$ preference for the updated, (input) heating-rate prescription. However, Bayesian model selection is often not practical outside of simulated data sets where we can control all aspects of data generation.
@@
While the fits with both prescriptions look good and do not raise suspicion, the same is not true of the parameter recovery. In Fig. 4, we show violin plots of the posteriors of the estimated ejecta masses and velocities of the two components, with the fit with the updated heating-rate prescription in blue and the older prescription in orange. The black crosses indicate the true input parameters. As illustrated, the incorrect prescription, while providing good fits by eye to the observations, fail to recover the true input parameters; estimating an ejecta mass of ${0.003}_{-0.0006}^{+0.0004}$ and ${0.007}\pm 0.002~{}M_{\odot}$ ($68\%$ credible interval) for the two components respectively, fully excluding the true input parameters of $0.015$ and $0.03~{}M_{\odot}$, respectively. Meanwhile, the velocities, while more consistent, also miss the true input velocities, and are more unconstrained by the older prescription relative to the updated heating-rate prescription, with the first component’s $68\%$ credible-interval width on the velocity of $18\%$ relative to $4\%$ with the updated prescription. The unconstrained posterior is a consequence of the older prescription providing no diagnostic power towards the velocity from the shape of the lightcurve, especially for the high electron fraction first component.
@@ 4 Uncertainty in heating rate prescriptions
For example, for low electron fractions, $Y_{e}\lesssim 0.2$, the choice between the Duflo-Zuker (Duflo & Zuker, 1995) and the Finite Range Droplet Model (FRDM) mass formulae (Möller et al., 1995) produce a difference of up to an order of magnitude in the heating rate (Barnes et al., 2016; Rosswog et al., 2017). A similar deviation is also seen with the Gogny-Hartree-Fock-Bogoliubov (HFB) (Goriely et al., 2009) mass model (Zhu et al., 2021; Barnes et al., 2021).
To investigate the effect of this uncertainty, we simulate synthetic kilonova observations, with both components with low electron fractions of $0.1$ and $0.15$, respectively, consistent with the ejecta of a neutron-star black hole merger. We fit these observations but increase the heating rate for electron fractions $Y_{e}\lesssim 0.2$ by a factor of $10$, with no change to the heating rate for other electron fractions. All input parameters apart from the electrons fractions are kept the same as the simulation in Sec. 3. The impact of a discrepancy in heating rate up to a factor of $10$ is severe. When fitting with $Y_{e}$ (or opacity) fixed, as commonly done in analyses for AT2017gfo (e.g., Coughlin et al., 2019; Nicholl et al., 2021), we fail to recover the correct input masses and velocities, with no posterior overlap with the true input. In particular, as the heating rate is assumed to be larger (i.e., more energy released per gram of material), we infer significantly smaller ejecta masses by up to a factor of 10. The large difference in heating rate also translates into a visually discernible and significantly worse quality of fit, partly as even with broad priors, there is no suitable combination of ejecta mass, velocity, and opacity compatible with the peak time of the simulated data. However, if we perform our analysis and keep the electron fractions as free parameters when fitting with a broad prior (we use a uniform prior from $Y_{e}=0.1-0.5$), we infer broad posteriors on ejecta masses consistent with the true input, but ejecta velocities and electron fraction posteriors with biases significantly larger than the statistical uncertainty. We note that even when fitting with electron fractions kept free, our fits are worse than with the correct input heating rate but not at a level discernible without robust statistical measures.
For studies where electron fractions could be fixed, due to e.g., spectral information, the large difference in fit quality suggest that data alone could distinguish between the heating-rate predictions for the different mass formulae. We note that if the effect of the mass formula is smaller, e.g., up to a factor of $\approx 3$, then photometric data alone is not sufficient to distinguish between the two mass formulae, with no discernible statistical difference in the quality of the fit. Analyses where electron fractions and therefore opacities can not be reliably constrained from e.g., the spectrum, and must be kept free when fitting, there is no discernible difference in fit quality. In such scenarios the data will lead to imprecise and inaccurate inferences on ejecta properties. In general, we expect distinguishing through data is likely to be difficult in practice due to weak constraints on the electron fraction from the spectrum as well as obfuscation from other modelling extensions or the effects of other kilonova modelling systematics discussed in Sec. 2. The discrepancies in inferred properties highlighted above are concerning, as if the discrepancy in the heating-rate due to the different mass formulae is not distinguishable through data alone, then inferred estimates of ejecta properties for lanthanide-rich kilonova ejecta are likely systematically biased by $\gtrsim 50\%$ for the ejecta velocities, while for ejecta masses the statistical uncertainties are underestimating the true uncertainty by up to an order of magnitude. Such uncertainties are far larger than the precision required for any precision constraints on the chemical enrichment of the Universe, or multi-messenger constraints on the equation of state or Hubble constant, while the bias on ejecta velocity, is even more problematic for accurate inferences due to nature of how bias grows when combining events in hierarchical analyses (Mandel et al., 2019).
Other sources of uncertainty in the effective heating rate, such as due to the nuclear network, reaction rates and thermalization are harder to quantify. However, given the range of estimates from multiple groups, we can conservatively estimate the uncertainty in the effective heating rate to be at least an order of magnitude the (Barnes et al., 2016; Hotokezaka & Nakar, 2020; Zhu et al., 2021; Barnes et al., 2021). This suggests that the uncertainties above are likely also valid for lanthanide-poor ejecta, i.e., the systematic uncertainty on the ejecta masses could be as high as an order of magnitude, while the ejecta velocities may also be systematically biased.
@@ 5 Conclusions
In this Letter, we have highlighted the dangers of using kilonova heating-rate prescription that are based on a very narrow $Y_{e}$ range such as the one from Korobkin et al. (2012). Although this prescription has been widely used as a key ingredient for estimating kilonova properties and designing survey strategies, it is unsuitable for electron fractions significantly different from $Y_{e}=0.04$, with substantial deviations in output for high electron fractions and at late times. Here, we have shown the differences in individual lightcurves and populations from an updated prescription following Rosswog & Korobkin (2024). We make these lightcurve models and heating-rate prescription publicly available in open-source package, Redback (Sarin et al., 2023) [^2] Redback is available at https://github.com/nikhil-sarin/redback.
@@
We have also shown the dangers of using the incorrect prescription by fitting realistic, simulated observations with the Vera Rubin Observatory. We find that while fits look convincing by eye and do not raise suspicion, they fail to recover the correct input ejecta properties with biases greater than the estimated statistical uncertainty. This has severe consequences for judging the relevance of neutron star mergers for the chemical enrichment of the cosmos, the interpretation of kilonovae and their use in multi-messenger studies to constrain the equation of state or Hubble constant. In a real observation, where we do not know the true ‘input’ parameters, this highlights the need to scrutinize estimated parameters with physical intuition and consistency with different theoretical studies due to the limitations inherent in detailed modelling of kilonovae.
Our study demonstrates that given current estimates of the uncertainty in effective or ‘naked’ heating rates, we can expect significant biases in the inferred ejecta properties. In particular, we find that analyses where the opacities (or electron fractions) are fixed fail to recover any input parameter, with biases significantly larger than the statistical uncertainty expected from near future observations. For analyses where electron fractions are not fixed, the ejecta masses are poorly measured but consistent with input parameters. However, the ejecta velocities and opacities in this scenario are significantly biased. Such imprecise mass measurements alongside biased opacities (or electron fractions) and velocities may pose significant challenges for correct interpretation of the data.
@@
We make a number of simplifying assumptions in our analysis in terms of models, including the treatment of thermalization efficiency and simplified relation of electron fraction to opacity. Moreover, our analysis was limited to a small subset of simulated scenarios. However, our limited study highlights the significant systematic uncertainties in inferring kilonovae properties from observations. This systematic uncertainty limits the precision of constraints on equation of state and Hubble constant from kilonovae, and should ideally be marginalised over when fitting observations such as AT2017gfo, or as additional components in gamma-ray burst afterglows, e.g., GRB 211211A and GRB230307A to ensure that the interpretation of these events are not affected. Moreover, this raises the need for caution regarding precision constraints based on kilonovae properties and interpretations of their role in the chemical enrichment of the Universe.
In this work, we have also introduced and showcased a framework and tool to better understand the impact of different modelling assumptions on inferences of parameters. In the future, we will extend this to work to capture different simulated scenarios and impacts of various other kilonovae modelling uncertainties and assumptions.
@@ 6 Acknowledgments
We are grateful to the anonymous referee for a detailed report with many helpful comments. We want to further thank Oleg Korobkin for his comments on an earlier version. N.S. is supported by a Nordita Fellowship. Nordita is supported in part by NordForsk. N.S. and S.R. acknowledge support from the Knut and Alice Wallenberg foundation through the “Gravity Meets Light” project (PIs: Rosswog & Jerkstrand). S.R. has been supported by the Swedish Research Council (VR) under grant number 2020-05044, by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012, by the Knut and Alice Wallenberg Foundation under grant Dnr. KAW 2019.0112, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and by the European Research Council (ERC) Advanced Grant INSPIRATION under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101053985).